OpenStudy (anonymous):

\[\Large \log^2x^3-20\log\sqrt x +1=? \] How many solutions does equation have? 1 2 3 4

OpenStudy (anonymous):

I can't start it :(

OpenStudy (anonymous):

I know that \[\Large \log^2x =2\log x\] and \[\Large \log_{x^t}a=\frac 1t \log_xa\] and something else but as much as I see I don't need the second one here ... :S Any hint??

OpenStudy (mertsj):

What is it equal to? If it is an equation, it must be equal to something.

OpenStudy (anonymous):

sorry... equal to 0 :$ my bad

OpenStudy (anonymous):

I thought that I write that :$

OpenStudy (mertsj):

\[\log_{10}x^2=2\log_{10}x \]

OpenStudy (mertsj):

\[(\log_{10}x )^{2}\neq2\log_{10}x \]

OpenStudy (anonymous):

\[\Large \log^2x =\log x\cdot \log x ??\]

OpenStudy (mertsj):

Yes

OpenStudy (turingtest):

\[\Large \log^2x^3-20\log\sqrt x +1=9\log^2x-10\log x+1\]it's not equal to 0 is it?cause then it would be quadratic in log x

OpenStudy (turingtest):

\[\Large \log^2x^3-20\log\sqrt x +1=0\]would make this problem a little more direct

OpenStudy (anonymous):

Is given: \[\Large \log^2x^3-20\log\sqrt x+1=0\]

OpenStudy (turingtest):

sweet, then you can use the quadratic formula from where I got you too

OpenStudy (turingtest):

\[\log^2x^3-20\log\sqrt x +1=9\log^2x-10\log x+1=0\]substitute\[\log x=y\]

OpenStudy (anonymous):

what about \[\Large \log \sqrt x\]

OpenStudy (anonymous):

Ufff.. I see sorry Let me try.. I'll see what I can do ;)

OpenStudy (mertsj):

Turing, sorry to be so slow, but how is log^2x^3=9log^2x

OpenStudy (turingtest):

not at all :)\[(\log x^3)^2=(3\log x)^2=9\log x\]

OpenStudy (mertsj):

Oh. I got it. Sorry.

OpenStudy (turingtest):

forgot the log^2 in the last part...

OpenStudy (mertsj):

Now it factors!!!

OpenStudy (turingtest):

sweet jeezus we might be able to solve it :D

OpenStudy (anonymous):

So I have... \[\Large x_{1/2}\frac{10\pm\sqrt{100-36}}{18}\] And \[\Large x_{1/2}\frac{10\pm 8}{18}\longrightarrow x_1=1 \quad ,x_2=\frac19\] Right?

OpenStudy (turingtest):

I don't know, I don't feel like checking mertsj says it factors. let's see what he gets

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

not factorable? ok let me see what I get...

OpenStudy (mertsj):

\[(9logx-1)(logx-1)=0\]

OpenStudy (turingtest):

I though it factored :) so it looks like we have the right answer... (for the log part at least)

OpenStudy (turingtest):

I suppose to finish this out we need to know what base the log is is it 10?

OpenStudy (anonymous):

Yes .. because there's no base given ...

OpenStudy (mertsj):

So 10^(1/9) and 10

OpenStudy (turingtest):

\[\log x=\frac19\implies x=\sqrt[9]10\]ew... ugly number well, ok not that ugly, but still

OpenStudy (turingtest):

at least the 10 is pretty :)

OpenStudy (anonymous):

LOL ... thanks guys so much, thanks in advance

OpenStudy (mertsj):

Very much so. Do these answers have to be checked in the original?

OpenStudy (mertsj):

Thank Turing. He's the brains of the outfit.

OpenStudy (turingtest):

Welcome, thanks for the compliment @Kreshnik why do you say "in advance" ? you assume I am going to keep helping that way? lol

OpenStudy (turingtest):

fyi you usually say "thanks in advance" before the person has actually done the favor

OpenStudy (turingtest):

(English lesson, lol)

OpenStudy (mertsj):

Oh. It only wanted to know how many solutions there are.

OpenStudy (mertsj):

A true liberal arts scholar!!

OpenStudy (turingtest):

haha^ screw that "how may" business, once you recognize it's quadratic in logx you know there will be 2 solutions, so I guess we could have stopped there

OpenStudy (turingtest):

...so long as the discriminant is not zero

OpenStudy (anonymous):

LOL ... hahaha , Since my own language it's not English , I've seen in a forum a guy who said that: Thanks in advance .... uhauhuha LOL INDEED he said that before the answer it was given, sorry if I made any mistake , I'm trying my best, I'm 18 years old, and I'm from Kosovo but I know English too (at least I think I do ! LOL ) hahaha.. @TuringTest thanks for English lessons :) ... can't stop laughing !

OpenStudy (turingtest):

That's cool, I'm learning Spanish and I teach English, so I know your pain

OpenStudy (mertsj):

Or one of the answers is not negative.

OpenStudy (turingtest):

good point @Mertsj

OpenStudy (mertsj):

You teach English??? And you're such a math whiz. Do you teach math too?

OpenStudy (anonymous):

I know spanish too... (I understand pretty much !) but I know English better !! ... (I don't know to write Spanish at all !)

OpenStudy (turingtest):

I am a tutor in English and math for the school where I am studying Spanish here in Mexico I actually suck at grammatical rules and such, but I get by being a native speaker.

OpenStudy (mertsj):

Be safe. And don't get involved in the drug wars!!

OpenStudy (turingtest):

No worries, that's a border problem. I'm in Guadalajara, to the south-central. Thanks though, and take care yourself :)

OpenStudy (mertsj):

Thanks. Will do.